In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate receive?
winner:___votes
loser:____votes

Answer by sprolden(40)

In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate receive?
winner:___votes
loser:____votes
1- Let x = the loser's votes.
2- The winning candidate has 220 more votes than the loser x+220
3- The winning candidate is represented by x+220 and the losing candidate is represented by x. 810 votes were cast in all x+(x+220)=810
4- Solve the equation for x by combining like terms
x+(x+220)=810
2x+220=810
5-Subtract 220 from each side
2x+220-220=810-220
2x=590
6- Divide each side by 2
2x/2=590/2
x=295
7- Earlier, I let x+220 be the winner's votes. The winner had 295+220 or 515 votes. The loser had x or 295 votes.
winner: 515 votes
loser: 295 votes
515+295=810 votes in all